TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Create new blog entry for Point 8


Created by Main.AntonioVergara on 2007-01-30 - 14:19

RAT in Point 8 - Tuesday 30/01

PDF RAT_20070130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x30x14x18/RAT_20070130.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-29 - 17:06

RAT in Point 8 - Monday Jan 29

PDF RAT_20070129_withAttachments.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x29x17x02/RAT_20070129_withAttachments.pdf

  • The short to ground in Sector 78 disappeared last Saturday
  • The cooling capacity was restored to 80% today
  • The mobile camera run will be planned tomorrow from Point 7
  • Preparation for Powering Tests will take place from tomorrow in UA83

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-26 - 12:27

RAT in point 8 - 26th January 2007

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x26x12x26/RAT_20070126.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 18:49

Preparation for the Powering Tests Sector 7-8

PDF RAT_78_070124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x18x48/RAT_78_070124.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 11:47

Preparation for Cool-Down Sectors 7-8,4-5,8-1

PDF RAT_20070125_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x11x45/RAT_20070125_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2007-01-24 - 16:56

la mise en froid se deroule tres normalement...

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-23 - 14:31

RAT in point 8 - 23th January 2007

PDF RAT_20070123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x23x14x25/RAT_20070123.pdf

Comments:

  • NEXT RAT will be on thursday 25th january 2007 at 8:30 -- Main.BorisBellesia - 23 Jan 2007 - 14:31

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-22 - 17:17

RAT in Point 8 - 22nd January 2007

PDF RAT_20070122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x22x17x13/RAT_20070122.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-19 - 19:37

RAT in Point 8 - Minutes - 19th Jan 2007

PDF RAT_20070119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x19x19x35/RAT_20070119.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-18 - 11:14

Cool-Down follow up - Jan 18th - Plots

PDF ELQA-HVQ_cooldown_2007-01-18.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/ELQA-HVQ_cooldown_2007-01-18.pdf

PDF Cooling_LHC_sector78_18Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/Cooling_LHC_sector78_18Jan07.pdf

Comments:

  • Inlet set temperature = 185K
  • Cooling Power = 300kW = 50% of nominal
  • Leakage Current Monitoring = ok

MircoCoccoli - 18 Jan 2007 - 10:51

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 17:24

RAT in point 8 - 17th January - Minutes

PDF RAT_20070117_with_NC_Mag_Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x17x17/RAT_20070117_with_NC_Mag_Planning.pdf

> ATTENTION: next RAT in point 8 on FRIDAY 19th at 8.30am <

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 16:43

Minutes of the meeting on Point 8 Warm Magnets Commissioning - 15th Jan

PDF Minutes_WarmMag_LSS8_15Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x15x56/Minutes_WarmMag_LSS8_15Jan07.pdf

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-16 - 18:33

RAT in point 8 - 16th January - Minutes

PDF RAT_20070116.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x16x18x25/RAT_20070116.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 16.01.07 Vac-Cryo Interlock tests – window of 1 day needed (P.Cruikshank-F.Millet)
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-01-16 - 10:17

First Cooldown QRL C to D

Images obtained from Francois Millet

PNG Cooldown.png:
Cooldown.png
First cooldown QRL 78
PNG First_cooldown_Q9R7.png:
First_cooldown_Q9R7.png
First cooldown Q9R7

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-15 - 18:42

RAT Meeting in Point 8 - Jan 15th

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x15x18x38/RAT_20070115.pdf

MircoCoccoli - 15 Jan 2007 - 18:40

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-12 - 18:46

RAT Meeting in Point 8 - Jan 12th

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x12x18x43/RAT_20070112.pdf

MircoCoccoli - 12 Jan 2007 - 18:45

Open Issues (Sect 78):

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 10.01.07 Connection of the NC Magnets in LSS8R (J-C.Guillaume)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-11 - 18:27

RAT Meeting P8 - 11th Jan Minutes

PDF RAT_20070111.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x11x18x26/RAT_20070111.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-10 - 15:13

RAT Meeting P8 - 10th Jan Minutes

PDF RAT_20070110.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x10x15x12/RAT_20070110.pdf

Comments:

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2007-01-09 - 11:05

Labels of the Water cooled cables following the ECR- EDMS 804188

Some pictures of the water-cooled cables with the labels showing they are following the ECR- EDMS 804188, when it's necessary

JPEG P1000209.JPG:
P1000209.JPG

JPEG P1000206.JPG:
P1000206.JPG

JPEG P1000207.JPG:
P1000207.JPG

JPEG P1000208.JPG:
P1000208.JPG

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-08 - 20:35

RAT Point 8 Minutes

PDF RAT_20070108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x08x20x33/RAT_20070108.pdf

RAT regarding Cool Down of Sect 78 and preparation for Cool Down of sectors 45 and 81

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-01-08 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login